# AgileDART: An Agile and Scalable Edge Stream Processing Engine

### v1

@@ Abstract
Edge applications generate a large influx of sensor data at massive scales. Under many time-critical scenarios, these massive data streams must be processed in a very short time to derive actionable intelligence. However, traditional data processing systems (e.g., stream processing systems, cloud-based IoT data processing systems) are not well-suited for these edge applications. This is because they often do not scale well with a large number of concurrent stream queries, do not support low-latency processing under limited edge computing resources, and do not adapt to the level of heterogeneity and dynamicity commonly present in edge computing environments. These gaps suggest a need for a new edge stream processing system that advances the stream processing paradigm to achieve efficiency and flexibility under the constraints presented by edge computing architectures.
We present AgileDart, an agile and scalable edge stream processing engine that enables fast stream processing of a large number of concurrently running low-latency edge applications’ queries at scale in dynamic, heterogeneous edge environments. The novelty of our work lies in a dynamic dataflow abstraction that leverages distributed hash table (DHT) based peer-to-peer (P2P) overlay networks to automatically place, chain, and scale stream operators to reduce query latencies, adapt to workload variations, and recover from failures; and a bandit-based path planning model that can re-plan the data shuffling paths to adapt to unreliable and heterogeneous edge networks. We show analytically and empirically that AgileDart outperforms Storm and EdgeWise on query latency and significantly improves scalability and adaptability when processing a large number of real-world edge stream applications’ queries.
@@ I Introduction
To facilitate fast stream processing for a vast number of concurrently running time-critical edge applications’ jobs at scale in a dynamic (varying workloads and bandwidths) and heterogeneous (different network layouts) edge environment, this paper proposes AgileDART, a next-generation scalable and adaptive edge stream processing engine. The key innovation is that AgileDART re-architects the stream processing system design. In sharp contrast to existing stream processing systems, there is no monolithic master. Instead, AgileDART involves all peer nodes to participate in operator placement, dataflow path shuffling, and operator scaling, thereby revolutionarily improving scalability and adaptivity.
@@
First, we highlight the challenges we face during the design of an efficient edge stream processing engine (Section II).
Second, we study the edge stream processing query model, the software architecture of existing stream processing systems, and discuss their limitations in supporting dynamic and heterogeneous edge environments. To the best of our knowledge, we are the first to observe the lack of scalability and adaptivity in stream processing systems for handling a large number of edge applications in practical edge settings (Section III).
Third, we design a novel dynamic dataflow abstraction to automatically place, chain and parallelize stream operators using the distributed hash table (DHT) based peer-to-peer (P2P) overlay networks. The main advantage of a DHT is that it avoids the original monolithic master. All peer nodes jointly make operator-mapping decisions. Nodes can be added or removed with minimal work around redistributing keys. This design allows our system to scale to extremely large numbers of nodes, applications, and jobs. To the best of our knowledge, we are the first to explore DHT-based P2P technologies to pursue extreme scalability in edge stream processing systems (Section IV).
Fourth, we introduce a novel bandit-based exploitation-exploration path planning model that can dynamically replan the stream data shuffling paths to adapt to the unreliable, heterogeneous edge networks (Section V).
Fifth, using DHTs, we decompose the stream processing system architecture from $1$:$n$ to $m$:$n$, which removes the centralized master and ensures that each edge zone can have an independent master for handling applications and operating autonomously without any centralized state. As a result of our distributed management, AgileDART improves overall query latencies for concurrently running applications and significantly reduces application deployment times. To the best of our knowledge, we offer the first stream processing engine to make it feasible to operate edge applications in a DevOps fashion (Section VI).
@@ IV-C Elastic Scaling Mechanism
We design a heuristic approach that adapts execution based on various factors: bottleneck (compute or bandwidth), operator (stateless or stateful), dynamics (long-term or short-term), dataflow structures, and noises. If there are computational bottlenecks, we scale up the problematic operators. The intuition is that when data queuing increases, automatically adding more instances to the system will avoid the bottleneck. We leverage the Secant root-finding method [52] to automatically calculate the optimal instance number based on the current system’s health value. The policy is pluggable. Let $f(x)$ represent the health score based on the input rate and the queue size ($10<f(x)<$ 1, with 1 being the highest score). Let $x_{n}$ and $x_{n-1}$ be the number of instances during phases $p_{n}$ and $p_{n-1}$. Then the number of instances required for the next phase $p_{n+1}$ such that $f\cong 1$ can be given by:
$x_{n+1}=x_{n}+(1-f(x_{n}))\times\frac{x_{n}-x_{n-1}}{f(x_{n})-f(x_{n-1})}$ (1)
@@ V-B Bandit-Based Data Shuffling Algorithm
In summary, this algorithm can efficiently identify the optimal path for data shuffling in unreliable edge networks, taking into account link heterogeneity and minimizing regret in the path selection process.
@@ VI Implementation
The efficiency comes from several factors. First, all nodes in the system are equal peers with the same rights and duties. Each node may act as one application’s worker, another application’s worker, a zone’s scheduler, or any combination of the above, resulting in all loads being evenly distributed. Second, the scheduler is no longer a bottleneck. Third, the system automatically creates more schedulers for application-intensive zones and fewer ones for sparse zones, thus scaling to extremely large numbers of nodes, applications, and queries.
@@ VII-C Scalability Analysis
We now show scalability: AgileDart decomposes the traditional centralized architecture of stream processing engines into a novel decentralized architecture for operator mapping and query scheduling. This structural change notably enhances the system’s capacity to scale effectively with a large number of concurrently running applications, application operators, and edge zones.
Distribution of operators. Figure 10(a) shows the mappings of AgileDart’s operators on edge nodes for 250, 500, 750, and 1,000 concurrently running applications. These applications run a mix of topologies with varying numbers of operators (with an average of 10). Figure 10(b) shows the normal probability plot of the number of operators per node. The results show that when deploying 250 and 500 applications, around 96.52% nodes host fewer than 3 operators; and when deploying 750 and 1000 applications, around 99.84% nodes host fewer than 4 operators. From Figure 10(a) and Figure 10(b), we can see that these applications’ operators are evenly distributed across all edge nodes. This is because AgileDart essentially leverages the DHT routing to map operators on edge nodes. Since the application’s dataflow topologies are different, their routing paths and the rendezvous points will also be different, resulting in operators well balanced across all edge nodes.
@@ VII-E Elastic Scaling Analysis
Health score. Figure 12(c) shows how the health score changes corresponding to Figure 12(a). Figure 12(d) shows how the health score changes corresponding to Figure 12(b). It’s worth noting that in cases where the objective of achieving a higher health score conflicts with the aim of improving throughput, it becomes necessary to strike a balance between the health score and system throughput. This balance can be achieved by adjusting the health score function to target a lower score.
@@ VII-F Data Shuffling Analysis
Path planning in real-world road map. Figure 13(b) presents the paths planned by AgileDart and its counterparts (end-to-end routing [81] and next-hop routing [82]) for an application after 50 trials (See Appendix B.1 and Appendix B.2 for more details of end-to-end routing and next-hop routing, respectively). In particular, end-to-end routing selects paths by leveraging the lower confidence bound (LCB) [83]; next-hop routing selects the next hop based on empirical packet delays; optimal routing selects the optimal path all the time. We considered an IoT-based application that needs to send 100 KB/s keyframes in the video data [84] from the source to the sink. We referred to [85, 86] to simulate network bandwidth over the road network. The expected packet delay of each link falls in the range from 50 ms to 250 ms, depending on the allocated bandwidth. The smaller the allocated bandwidth of a link, the longer its expected packet delay. The results show that optimal routing can find the shortest path with 5.9 second of packet delay, and the packet delay of the path planned by AgileDart is 7.8 second, which is closer to optimal routing than end-to-end routing and next-hop routing.
@@
Number of trials. Figure 16 shows the packet delays of paths selected by different algorithms at the different number of trials. We can see that AgileDart finds the optimal path at the $26_{th}$ trial, which is faster than the $33_{rd}$ trial and the $38_{th}$ trial of next-hop routing and end-to-end routing, respectively. Moreover, AgileDart keeps trying different paths despite finding the optimal one. This is because the exploration factor guides AgileDart to explore some paths worth trying.
@@ VIII Related Work
Category 3: Edge-based data processing systems. In such a model, data processing is performed at the edge without connectivity to a cloud backend [11, 12, 13]. This requires installing a hub device at the edge to collect data from other edge devices and perform data processing. These solutions, however, are limited by the computational capabilities of the hub service and cannot support distributed data-parallel processing across many devices and thus have limited throughput. It may also introduce a single point of failure once the hub device fails.
Category 4: Serverless-based stream processing systems. In such a model, data processing is performed on serverless platforms designed to handle event-driven workloads in a scalable and cost-effective manner, benefiting from the automatic scaling capabilities of serverless architectures that dynamically respond to demand fluctuations without manual intervention [91, 92, 93, 94]. This model, however, may inherit cold start latency in serverless functions, arising from function initialization, which presents a critical challenge to real-time responsiveness—an essential factor for time-sensitive edge stream applications where near-instantaneous processing is imperative.
@@
To the best of our knowledge, Edgent [9], EdgeWise [10], and Frontier [33] are the only other stream processing engines tailored for the edge. They all point out the criticality of edge stream processing, but no effective solutions were proposed towards scalable and adaptive edge stream processing. Edgent [9] is designed for data processing at individual edge devices rather than full-fledged distributed stream processing. EdgeWise [10] develops a congestion-aware scheduler to reduce backpressure, but it can not scale well due to the centralized bottleneck. Frontier [33] develops replicated dataflow graphs for fault-tolerance, but it ignores the edge dynamics and heterogeneity.

### v2

@@ Abstract
Edge applications generate a large influx of sensor data on massive scales, and these massive data streams must be processed shortly to derive actionable intelligence. However, traditional data processing systems are not well-suited for these edge applications as they often do not scale well with a large number of concurrent stream queries, do not support low-latency processing under limited edge computing resources, and do not adapt to the level of heterogeneity and dynamicity commonly present in edge computing environments. As such, we present AgileDart, an agile and scalable edge stream processing engine that enables fast stream processing of many concurrently running low-latency edge applications’ queries at scale in dynamic, heterogeneous edge environments. The novelty of our work lies in a dynamic dataflow abstraction that leverages distributed hash table-based peer-to-peer overlay networks to autonomously place, chain, and scale stream operators to reduce query latencies, adapt to workload variations, and recover from failures and a bandit-based path planning model that re-plans the data shuffling paths to adapt to unreliable and heterogeneous edge networks. We show that AgileDart outperforms Storm and EdgeWise on query latency and significantly improves scalability and adaptability when processing many real-world edge stream applications’ queries.
@@ I Introduction
To facilitate fast stream processing for a vast number of concurrently running time-critical edge applications’ jobs at scale in a dynamic (varying workloads and bandwidths) and heterogeneous (different network layouts) edge environment, this paper proposes AgileDart, a next-generation scalable and adaptive edge stream processing engine. In sharp contrast to existing stream processing systems relying on a single monolithic master, the key innovation of AgileDart lies in a decentralized architecture in which all peer nodes autonomously place operators, plan dataflow paths for data shuffling, and scale operators, thereby dramatically improving scalability and adaptivity.
@@
First, we highlight the challenges when designing an efficient edge stream processing engine (Section II) and then study existing stream processing systems’ query model and software architecture and their limitations in supporting dynamic and heterogeneous edge environments. To the best of our knowledge, we are the first to observe the lack of scalability and adaptivity in existing stream processing systems for handling massive edge applications in practical edge settings (Section III).
Second, we design a novel dynamic dataflow abstraction that autonomously places, chains, and parallelizes stream operators using the distributed hash table (DHT) based peer-to-peer (P2P) overlay networks. The main advantage of using DHT is that it avoids the original monolithic master, and P2P overlay networks enable all peer nodes to make operator-mapping decisions jointly. Moreover, nodes can join or leave the system with minimal work by redistributing keys. Hence, the design allows our system to scale to vast numbers of nodes, applications, and jobs simultaneously. To the best of our knowledge, we are the first to explore DHT-based P2P technologies to pursue extreme scalability in edge stream processing systems (Section IV).
Third, we introduce a novel bandit-based exploitation-exploration path planning model to reinforce the DHT-based routing protocols in the dynamic dataflow abstraction. Instead of relying only on the DHT-based routing tables with fixed prefixes, the model dynamically replans the stream data shuffling paths to adapt to unreliable and heterogeneous edge networks (Section V).
Fourth, using DHTs, we decompose the stream processing system architecture from one master to $N$ workers to $M$ masters to $N$ workers, which removes a monolithic centralized master and ensures that each edge zone can have an independent master for handling applications and operating individually without any centralized state. As a result of our distributed management, AgileDart improves overall query latencies for concurrently running applications and significantly reduces application deployment times. To the best of our knowledge, we offer the first stream processing engine to enable operating edge applications in a DevOps fashion (Section VI).
@@ IV-C Elastic Scaling Mechanism
We design a heuristic approach that adapts execution based on various factors: bottleneck (compute or bandwidth), operator (stateless or stateful), dynamics (long-term or short-term), dataflow structures, and noises. If there are computational bottlenecks, we scale up the problematic operators. The intuition is that when data queuing increases, automatically adding more instances to the system will avoid the bottleneck. We leverage the Secant root-finding method [52] to automatically calculate the optimal instance number based on the current system’s health value. The policy is pluggable. Suppose each instance can increase input rate and queue size by $r$ and $q$, respectively. In phase $p_{n}$, the rate at which new tasks or requests arrive at the system is $R_{p_{n}}$, and the number of tasks waiting to be processed is $Q_{p_{n}}$. Let $x_{n}$ and $x_{n-1}$ be the number of instances during phases $p_{n}$ and $p_{n-1}$. We define health score $f(x_{n})$ as
$f(x_{n})=\alpha\frac{x_{n}\times r}{R_{p_{n}}}+(1-\alpha)\frac{x_{n}\times q}{Q_{p_{n}}},$
where $\alpha$ is the weight tuning the importance of two terms. Then the number of instances required for the next phase $p_{n+1}$ such that $f\cong 1$ can be given by:
$x_{n+1}=x_{n}+(1-f(x_{n}))\times\frac{x_{n}-x_{n-1}}{f(x_{n})-f(x_{n-1})}.$
@@ V-B Bandit-Based Data Shuffling Algorithm
In summary, this algorithm can efficiently identify the optimal path for data shuffling in unreliable edge networks, taking into account link heterogeneity and minimizing regret in the path selection process. We conduct complexity analysis in Appendix B.
@@ VII-C Scalability Analysis
We now show scalability: AgileDart decomposes the traditional centralized architecture of stream processing engines into a novel decentralized architecture for operator mapping and query scheduling. This structural change notably enhances the system’s capacity to scale effectively with massive concurrently running applications, application operators, and edge zones. We deploy operators from up to 1,000 applications across 10,000 edge nodes over 20 edge zones to study whether AgileDart achieves an even distribution of operators or overburdens some nodes.
Distribution of operators. Figure 10(a) shows the mappings of AgileDart’s operators on edge nodes for 250, 500, 750, and 1,000 concurrently running applications. These applications run a mix of topologies with varying numbers of operators (with an average of 10). Figure 10(b) shows the normal probability plot of the number of operators per node. The results show that when deploying 250 and 500 applications, around 97.85% nodes host fewer than 3 operators; and when deploying 750 and 1,000 applications, around 99.89% nodes host fewer than 4 operators. From Figure 10(a) and Figure 10(b), we can see that these applications’ operators are evenly distributed across all edge nodes. This is because AgileDart essentially leverages the DHT routing to map operators on edge nodes. Since the application’s dataflow topologies are different, their routing paths and the rendezvous points will also be different, resulting in operators well balanced across all edge nodes.
@@ VII-D Failure Recovery Analysis
Computation and communication overhead comparison of AgileDart, Replication, and Checkpointing is presented in Appendix C.
@@ VII-E Elastic Scaling Analysis
Health score. Figure 12(c) shows how the health score changes corresponding to Figure 12(a). Figure 12(d) shows how the health score changes corresponding to Figure 12(b). It’s worth noting that in cases where the objective of achieving a higher health score conflicts with the aim of improving throughput, it becomes necessary to strike a balance between the health score and system throughput. This balance can be achieved by adjusting the health score function to target a lower score. More analysis on health scores can be found in Appendix D.
@@ VII-F Data Shuffling Analysis
Path planning in real-world road map. Figure 13(b) presents the paths planned by AgileDart and its counterparts (end-to-end routing [80] and next-hop routing [81]) for an application after 50 trials (See Appendix E.1 and Appendix E.2 for more details of end-to-end routing and next-hop routing, respectively). In particular, end-to-end routing selects paths by leveraging the lower confidence bound (LCB) [82]; next-hop routing selects the next hop based on empirical packet delays; optimal routing selects the optimal path all the time. We considered an IoT-based application that needs to send 100 KB/s keyframes in the video data [83] from the source to the sink. We referred to [84, 85] to simulate network bandwidth over the road network. The expected packet delay of each link falls in the range from 50 ms to 250 ms, depending on the allocated bandwidth. The smaller the allocated bandwidth of a link, the longer its expected packet delay. The results show that optimal routing can find the shortest path with 5.9 second of packet delay, and the packet delay of the path planned by AgileDart is 7.8 second, which is closer to optimal routing than end-to-end routing and next-hop routing.
@@
Number of trials. Figure 16 shows the packet delays of paths selected by different algorithms at the different number of trials. We can see that AgileDart finds the optimal path at the $26_{th}$ trial, which is faster than the $33_{rd}$ trial and the $38_{th}$ trial of next-hop routing and end-to-end routing, respectively. Moreover, AgileDart keeps trying different paths despite finding the optimal one. This is because the exploration factor guides AgileDart to explore some paths worth trying. We conduct more convergence analysis in Appendix F.
@@ VIII Related Work
Category 3: Edge-based data processing systems. In such a model, data processing is performed at the edge without connectivity to a cloud backend [11, 12, 13, 90, 91, 92, 93] or with cloud-edge cooperation [94, 95, 96]. This requires installing a hub device at the edge to collect data from other edge devices and perform data processing. GeoEkuiper [94] abstracts spatial data operations into necessary SQL statements functions and leverages cloud-native messaging to improve throughput in resource-constrained edge devices. Shepherd proposes efficient dynamic placement of operators over the hierarchical structure between edge and cloud. Amnis [91] aims at reducing end-to-end latency and overall throughput by considering data locality and resource constraints when generating physical DAG workflows. EQUALITY [92] proposes heuristic approaches to solving the trade-off between latency and overhead of data quality checks in edge stream processing. Beaver [93] aims to consider varying network bandwidths and heterogeneous computing resources for operator placement. These solutions, however, either require centralized coordination or are limited by the computational capabilities of the hub service and cannot support distributed data parallel processing across many devices, thus limiting throughput. It may also introduce a single point of failure once the hub device fails.
Category 4: Serverless-based stream processing systems. In such a model, data processing is performed simultaneously on serverless and serverful platforms. Serverless is leveraged to handle event-driven workloads in a scalable and cost-effective manner, benefiting from the automatic scaling capabilities of serverless architectures that dynamically respond to demand fluctuations without manual intervention [97, 98, 99, 100, 93, 101, 102, 103]. Sponge [97] leverages serverless framework instances to offload unpredictable increases in workloads from existing VMs with low latency and cost. Smartpick [103] exploits serverless and VMs together to realize composite benefits, i.e., agility from serverless and better performance with reduced cost from VMs. Storm-RTS [101] proposes translating stream operations into function-as-a-service invocations to achieve transparent and predictable performance. This model complements AgileDart by offering bursty workload offloading and transparent and predictable performance, whereas AgileDart provides scalable system support.
@@
Edgent [9], EdgeWise [10], and Frontier [33] are other stream processing engines tailored for the edge. They all point out the criticality of edge stream processing, but no effective solutions were proposed for scalable and adaptive edge stream processing. Edgent [9] is designed for data processing at individual edge devices rather than full-fledged distributed stream processing. EdgeWise [10] develops a congestion-aware scheduler to reduce backpressure, but it can not scale well due to the centralized bottleneck. Frontier [33] develops replicated dataflow graphs for fault-tolerance, but it ignores the edge dynamics and heterogeneity.